On July 21, 2019, Andrew purchased and placed in service a new car. The purchase price...

  1. On July 21, 2019, Andrew purchased and placed in service a new car. The purchase price was $52,000. This was the only business-use asset Andrew acquired in 2019. He used the car 85% of the time for business and 15% for personal use and maintains proper documentation of use. Andrew would like to deduct the maximum amount possible. Calculate the total deduction Andrew can take with respect to the car for 2019.
  2. Kendra Brown, a sole proprietor, acquires a new five-year asset on April 17, 2019, for $74,000 (for business-use). This is the only asset Kendra acquired during the year. Kendra does NOT elect §179 (and elects out of bonus depreciation). On September 1, 2020, she sells the asset.
  3. Juan acquired an apartment building on November 14, 2019, for $2.5 million. The value of the land was $325,000. Juan sold the apartment building and land on October 22, 2026.
    1. Determine his cost recovery for 2019
    2. Determine his cost recovery for 2026
    3. Determine his adjusted basis for the building as of the date of sale.
    4. If Juan sold the entire property for $2.7 million, how much was his gain?
    1. Determine the cost recovery for 2019
    2. Determine the cost recovery for 2020
Answer #1

A)

Determine his cost recovery for 2019

$ 2.5 mil – 325,000 = 2,175,000 * .00455 = 9,896.25

B)

Determine his cost recovery for 2026

$ 2,175,000 * .3636 = 79,083

C)

Determine his adjusted basis for the building as of the date of sale.

= $ 2,175,000 – 9,896.25 – 79,083 –[(2,175,000 *.3636) *6] = 1,611,522.75

D)

If Juan sold the building for $2.7 million, how much was his gain :

42,700,000 – 1,611,522.75 = 1,088,477.251

* Andrew can take with respect to the cost for 2019 = $52,000 * 85% = 44,200

* The cost recovery for kendra brown :

a. The cost recovery for 2019

= $74,000 * .20 = 14,800

b. The cost recovery for 2020

= $74,000 * .50 = 37,000 * .32 = 11,840